Solve for Y(s), the Laplace transform of the solution y(t) to the initial value problem below. y +6y=t? - 2, y(0) = 0, y(0
Answer #1

colle S3 Sol we have y+6y=+²-2 , 4(0) = 0 , 4(o)=-5 Take haplace both side: L(4) + 6 Lly) = (1+²) - 2 L(1) SPY05) - sy®) -